# A Non-Expert’s Introduction to Data Ethics for Mathematicians

### v4

@@ Abstract.
In parts of this chapter, I drew on materials from a class that Matt Salganik has taught on computational social science at Princeton University. Johan Ugander suggested several excellent resources and research studies. I thank Jacob Foster, Mark Handcock, Andrei Kleshin, Peter Mucha, Dennis Müller, Deanna Needell, Massimo Stella, Alexandria Volkening, Chris Wiggins, Richard Yeh, and an anonymous referee for helpful comments. Finally, I thank my fellow Short Course organizers (Heather Zinn Brooks, Michelle Feng, and Alexandria Volkening) for their continuing support on this project and elsewhere, their collaboration, and their friendship.
@@ 1. Introduction
The mathematical, statistical, and computational sciences are interconnected with human communities and society at large [cam, eth, MC23, Mos21]. Neither our education nor our research occur in a vacuum, and many of our algorithms and other methods are now applied to social systems [WSAO+21, CM23]. However, it is not traditional in mathematical-science education to discuss data ethics and other ethical considerations [RSMCM22, Mül22]. What we choose to teach (and choose not to teach) impacts what our mentees do with their education. Consequently, it is crucial for mathematical scientists (and others) to think carefully about ethics and engage with it throughout their careers [CM24]. See [LMP18] for a short book on ethics and data science, [FT16] for a theme journal issue on data ethics, [Vél21] for a wide-ranging handbook of digital ethics, and [MCF22] for a recent discussion of the ethics landscape in mathematics and advocacy of a “Hippocratic Oath” for mathematics.
@@ 2.1. The Importance of Data Ethics in the Mathematics Community and Mathematical-Science Curricula
Mathematical research also does not occur in a vacuum, especially when we apply our algorithms and other methods to social systems [CM23]. It is not a traditional part of curricular education or research education in the mathematical sciences to teach our students and the other junior community members about data ethics and other ethical considerations [RSMCM22, Mül22], but this needs to change. [^3] One positive development is that some computer-science conferences now require authors to include ethics statements in their submitted papers. What we choose to teach (and choose not to teach, or choose to mention in only a cursory way) impacts what our students and other mentees do with their education. Other disciplines (e.g., the social and medical sciences) have thought a lot more about ethics than mathematics (and allied disciplines), and we should be guided by the best practices that they have developed. Many of these best practices arose in the aftermath of ethically problematic studies, and ethical guidelines have developed as people have tried to learn from past mistakes. As in other arenas, the mathematics community has had major ethical problems (e.g., through various forms of social toxicity), but these problems traditionally have arisen in issues other than the scientific use and abuse of data. This situation has changed; mathematicians are now also facing data-related issues directly in their research. What we learn, teach, and do needs to catch up to this reality.
@@
It is important for mathematical scientists (and others) to think carefully about ethics and engage with it throughout their careers [CM24]. We need to incorporate data ethics into the core education in the mathematical, statistical, and computational sciences. We are increasingly using social, animal, and human data (including potentially personal data) in our research. We need to think carefully about when it is appropriate to use such data and when it is not appropriate, and there needs to be systematic training to help mathematical scientists confront these issues. See [LMP18] for a short book on ethics and data science, [FT16] for a theme journal issue on data ethics, [Vél21] for a wide-ranging handbook of digital ethics, [MCF22] for a discussion of the ethics landscape in mathematics and advocacy of a “Hippocratic Oath” for mathematics (also see [D’A22]), and [Sku21] for an article with a useful discussion and pointers to several relevant resources. See [Ass18] for the Code of Ethics and Professional Conduct of the Association for Computing Machinery (ACM). Buell et al. [BPT22] leveraged the ethical-practice standards of the ACM and the American Statistical Association (ASA), which both represent disciplines that are relevant to mathematics, in a survey of mathematicians about ethical standards. See [TPB24a, TPB24b] for further discussions of the guidelines that they developed for ethical mathematical practice, and see [WT24] for associated teaching materials. To help acknowledge issues like data ethics in the mathematical sciences, it is also relevant to include pertinent information on departmental websites. For an example website, which I helped produce at University of Oxford in the aftermath of a study [TMP12] by my collaborators and me that uses Facebook data, see [MI].
@@ 2.2. Course Material on Data Ethics and Society
Several courses have websites with much useful information about data ethics and related topics. I will highlight a few examples. Matthew Salganik (Department of Sociology, Princeton University) has taught a graduate course (from which I drew some material for the present chapter) called Computational Social Science: Social Research in the Digital Age. The material for the course’s Fall 2016 edition is available at [Sal16]. It includes material that appeared later in book form [Sal17]. Johan Ugander (Department of Management Science and Engineering, Stanford University) has taught a graduate course called Data Ethics and Privacy. See [Uga20] for the course website, from which I drew several articles that I mention in Section 6. Johan Ugander’s graduate course Social Algorithms [Uga23] also has excellent resources. Chris Bail (Department of Sociology, Duke University) has taught a graduate course called Data Science & Society. See [Bai22] for its course materials, schedule, and YouTube videos. A recent book by Chris Wiggins (Department of Applied Physics and Applied Mathematics, Columbia University) and Matthew L. Jones (Department of History, Columbia University) [WJ23] is an adaptation of their course Data: Past, Present, and Future [WW24].
@@ 4.3. Privacy Concerns and Practical Considerations for Open Data and Code
In Section 4.1, I mentioned that it is important to indicate how one has anonymized data in scholarly works. When is data actually “anonymous”, and is it ever possible to “fully” anonymize data? Consider the following scenario, which was discussed in [Sal16]. Suppose that we have a data set of medical records of individuals that includes their names, their home addresses, the zip codes of these addresses, their birth dates, their sexes, their ethnicities, the dates that each individual visited a doctor, the medical diagnoses, the medical procedures, and the prescribed medications. Now suppose that we “anonymize” this data set by removing the names and home addresses of all individuals. We now have a data set of “anonymized” medical records. Suppose that we obtain a data set of voting records and that this data set includes names, home addresses, political-party affiliations, voter-registration dates, zip codes, birth dates, and sexes. Both data sets include the zip codes, birth dates, and sexes of the individuals who are common to the two data sets. One can use such data — namely, data that is common to these two data sets — to de-anonymize people in the supposedly “anonymized” data set of medical records [NS08]. An infamous example of data de-anonymization by combining data sets led to the cancellation of the sequel to the Netflix Prize [Wikb].
@@ 5. Ethical Principles and Other Considerations in the Scientific Use of Human Data
In this section, I discuss some ethical principles and other considerations in the scientific use of human data. [^8] Parts of my discussion draw heavily from material in Matt Salganik’s course on computational social science [Sal16]. I also discuss Institutional Review Boards (IRBs). [^9] At some institutions, the backronym for IRB is Internal Review Board. Much of my discussion also applies to data from nonhuman animals, but I am focusing on human data in this chapter, so I typically phrase my exposition in human terms.
@@ 6.1. Algorithmic Biases and Other Biases
Biases in computer systems, machine learning, and AI go far beyond only algorithmic bias [FN96]. See [MMS+21, BHN23, CDNSG23] for reviews of bias and “fairness” in machine learning, and see [N+20] for an introductory survey of bias in data-driven AI systems. As advocated in [KA21], it is important to go beyond notions of mere algorithmic “fairness” (which focuses on intra-group versus inter-group differences). One must also analyze (1) inequality and the causal impact of algorithms and (2) the distribution of power. It is also important to be cognizant of traps that can beset work on “fair” machine learning in sociotechnical systems [SBF+19]. In “fair” machine learning, it is necessary to carefully examine context-specific consequences [CDNSG23].
@@ 7. Conclusion
In this chapter, I have discussed many aspects of data ethics in education, publishing, and research. In some of my discussions, I have also touched on data and justice. It is important to distinguish between ethics and justice, including when it comes to data [Kit14]. Ethics presupposes laws and social norms, and then one considers what is morally appropriate within those frameworks. Justice is concerned with how to change laws and modify social norms to attain broader equity. Both ethics and justice are core aspects of our interactions with data.

### v5

@@ Abstract.
In parts of this chapter, I drew on materials from a class that Matt Salganik has taught on computational social science at Princeton University. Johan Ugander suggested several excellent resources and research studies. I thank Jacob Foster, Mark Handcock, Andrei Kleshin, Peter Mucha, Dennis Müller, Deanna Needell, Matt Salganik, Massimo Stella, Alexandria Volkening, Chris Wiggins, Richard Yeh, and an anonymous referee for helpful comments. Finally, I thank my fellow Short Course organizers (Heather Zinn Brooks, Michelle Feng, and Alexandria Volkening) for their continuing support on this project and elsewhere, their collaboration, and their friendship.
@@ 1. Introduction
The mathematical, statistical, and computational sciences are interconnected with human communities and society at large [cam, eth, MC23, Mos21]. Neither our education nor our research occur in a vacuum, and many of our algorithms and other methods are now applied to social systems [WSAO+21, CM23]. However, it is not traditional in mathematical-science education to discuss data ethics and other ethical considerations [RSMCM24, Mü22]. What we choose to teach (and choose not to teach) impacts what our mentees do with their education. Consequently, it is crucial for mathematical scientists (and others) to think carefully about ethics and engage with it throughout their careers [CM24]. See [LMP18] for a short book on ethics and data science, [FT16] for a theme journal issue on data ethics, [Vé21] for a wide-ranging handbook of digital ethics, and [MCF22] for a recent discussion of the ethics landscape in mathematics and advocacy of a “Hippocratic Oath” for mathematics.
@@ 2.1. The Importance of Data Ethics in the Mathematics Community and Mathematical-Science Curricula
Mathematical research also does not occur in a vacuum, especially when we apply our algorithms and other methods to social systems [CM23]. It is not a traditional part of curricular education or research education in the mathematical sciences to teach our students and the other junior community members about data ethics and other ethical considerations [RSMCM24, Mü22], but this needs to change. [^3] One positive development is that some computer-science conferences now require authors to include ethics statements in their submitted papers. What we choose to teach (and choose not to teach, or choose to mention in only a cursory way) impacts what our students and other mentees do with their education. Other disciplines (e.g., the social and medical sciences) have thought a lot more about ethics than mathematics (and allied disciplines), and we should be guided by the best practices that they have developed. Many of these best practices arose in the aftermath of ethically problematic studies, and ethical guidelines have developed as people have tried to learn from past mistakes. As in other arenas, the mathematics community has had major ethical problems (e.g., through various forms of social toxicity), but these problems traditionally have arisen in issues other than the scientific use and abuse of data. This situation has changed; mathematicians are now also facing data-related issues directly in their research. What we learn, teach, and do needs to catch up to this reality.
@@
It is important for mathematical scientists (and others) to think carefully about ethics and engage with it throughout their careers [CM24]. We need to incorporate data ethics into the core education in the mathematical, statistical, and computational sciences. We are increasingly using social, animal, and human data (including potentially personal data) in our research. We need to think carefully about when it is appropriate to use such data and when it is not appropriate, and there needs to be systematic training to help mathematical scientists confront these issues. See [LMP18] for a short book on ethics and data science, [FT16] for a theme journal issue on data ethics, [Vé21] for a wide-ranging handbook of digital ethics, [MCF22] for a discussion of the ethics landscape in mathematics and advocacy of a “Hippocratic Oath” for mathematics (also see [D’A22]), and [Sku21] for an article with a useful discussion and pointers to several relevant resources. See [Ass18] for the Code of Ethics and Professional Conduct of the Association for Computing Machinery (ACM). Buell et al. [BPT22] leveraged the ethical-practice standards of the ACM and the American Statistical Association (ASA), which both represent disciplines that are relevant to mathematics, in a survey of mathematicians about ethical standards. See [TPB24a, TPB24b] for further discussions of the guidelines that they developed for ethical mathematical practice, and see [WT24] for associated teaching materials. To help acknowledge issues like data ethics in the mathematical sciences, it is also relevant to include pertinent information on departmental websites. For an example website, which I helped produce at University of Oxford in the aftermath of a study [TMP12] by my collaborators and me that uses Facebook data, see [MI].
@@ 2.2. Course Material on Data Ethics and Society
Several courses have websites with much useful information about data ethics and related topics. I will highlight a few examples. Matt Salganik (Department of Sociology, Princeton University) has taught a graduate course (from which I drew some material for the present chapter) called Computational Social Science: Social Research in the Digital Age. The material for the course’s Fall 2016 edition is available at [Sal16]. It includes material that appeared later in book form [Sal17]. Johan Ugander (Department of Management Science and Engineering, Stanford University) has taught a graduate course called Data Ethics and Privacy. See [Uga20] for the course website, from which I drew several articles that I mention in Section 6. Johan Ugander’s graduate course Social Algorithms [Uga23] also has excellent resources. Chris Bail (Department of Sociology, Duke University) has taught a graduate course called Data Science & Society. See [Bai22] for its course materials, schedule, and YouTube videos. A recent book by Chris Wiggins (Department of Applied Physics and Applied Mathematics, Columbia University) and Matthew L. Jones (Department of History, Columbia University) [WJ23] is an adaptation of their course Data: Past, Present, and Future [WW24].
@@ 4.3. Privacy Concerns and Practical Considerations for Open Data and Code
In Section 4.1, I mentioned that it is important to indicate how one has anonymized data in scholarly works. When is data actually “anonymous”, and is it ever possible to “fully” anonymize data? Consider the following scenario, which is discussed in [Sal16, Sal17]. Suppose that we have a data set of medical records of individuals that includes their names, their home addresses, the zip codes of these addresses, their birth dates, their sexes, their ethnicities, the dates that each individual visited a doctor, the medical diagnoses, the medical procedures, and the prescribed medications. Now suppose that we “anonymize” this data set by removing the names and home addresses of all individuals. We now have a data set of “anonymized” medical records. Suppose that we obtain a data set of voting records and that this data set includes names, home addresses, political-party affiliations, voter-registration dates, zip codes, birth dates, and sexes. Both data sets include the zip codes, birth dates, and sexes of the individuals who are common to the two data sets. One can use such data — namely, data that is common to these two data sets — to de-anonymize people in the supposedly “anonymized” data set of medical records [NS08]. An infamous example of data de-anonymization by combining data sets led to the cancellation of the sequel to the Netflix Prize [Wikb].
@@ 5. Ethical Principles and Other Considerations in the Scientific Use of Human Data
In this section, I discuss some ethical principles and other considerations in the scientific use of human data. [^8] Parts of my discussion draw heavily from material in Matt Salganik’s course on computational social science [Sal16]. See his associated book [Sal17] for further discussion. I also discuss Institutional Review Boards (IRBs). [^9] At some institutions, the backronym for IRB is Internal Review Board. Much of my discussion also applies to data from nonhuman animals, but I am focusing on human data in this chapter, so I typically phrase my exposition in human terms.
@@ 5.2. Four Key Principles
I now enumerate a few critical ethical principles, which are discussed in detail in chapter 6 of [Sal17].
@@ 6.1. Algorithmic Biases and Other Biases
Biases in computer systems, machine learning, and AI go far beyond only algorithmic bias [FN96, HV24]. See [MMS+21, BHN23, CDNSG23] for reviews of bias and “fairness” in machine learning, and see [N+20] for an introductory survey of bias in data-driven AI systems. As advocated in [KA21], it is important to go beyond notions of mere algorithmic “fairness” (which focuses on intra-group versus inter-group differences). One must also analyze (1) inequality and the causal impact of algorithms and (2) the distribution of power. It is also important to be cognizant of traps that can beset work on “fair” machine learning in sociotechnical systems [SBF+19]. In “fair” machine learning, it is necessary to carefully examine context-specific consequences [CDNSG23].
@@ 7. Conclusion
In this chapter, I have discussed many aspects of data ethics in education, publishing, and research. In some of my discussions, I have also touched on issues of data and justice. It is important to distinguish between ethics and justice, including when it comes to data [Kit14]. Ethics presupposes laws and social norms, and then one considers what is morally appropriate within those frameworks. Justice is concerned with how to change laws and modify social norms to attain broader equity. Both ethics and justice are core aspects of our interactions with data.